A 0.18 M solution of a weak base A‑ reacts with H2O such that at equilibrium 99.7% of the compound remains in the A- form. Given this information, what is the pH of this solution to the nearest tenths?
Answer #1

A- + H2O —> AH + OH-

99.7 % is remaining as A-

so,

[OH-] = 0.3 % of initial concentration

[OH-] = 0.3 % of 0.18 M

[OH-] = 0.3 * 0.18 / 100 = 5.4*10^-4 M

use:

pOH = -log [OH-]

= -log (5.4*10^-4)

= 3.3

use:

PH = 14 - pOH

= 14 - 3.3

= 10.7

Answer: 10.7